OpenStudy (jiteshmeghwal9):

Ramesh recieved Rs.32768 from a bank at the rate of \(12\frac{1}{2}\)% per annum calculated quarterly. Find the amount he will have to return after 9 months in order to clear his debts.

OpenStudy (jiteshmeghwal9):

@ganeshie8 sir @lgbasallote sir plz help :)

OpenStudy (jiteshmeghwal9):

@Zarkon plz help :)

OpenStudy (jiteshmeghwal9):

@lgbasallote sir plz help me :)

ganeshie8 (ganeshie8):

Rs. 32,768 @ 12.5 % compounded quarterly.

OpenStudy (jiteshmeghwal9):

yeah !

ganeshie8 (ganeshie8):

quarterly interest rate = 12.5/4 = 3.125 %

OpenStudy (jiteshmeghwal9):

yes !

ganeshie8 (ganeshie8):

every quarter, that is every 3 months, the bank will calculate interest on outstanding amount, and add the interest bill it to his account. so, every quarter his debt increases by the interest amount

ganeshie8 (ganeshie8):

3 quarters in 9 months

OpenStudy (jiteshmeghwal9):

k!

OpenStudy (jiteshmeghwal9):

Thanxx !

ganeshie8 (ganeshie8):

total amount he will have to return with interest = \(A (1 + r)^n\) = \(32,768(1 + 3.125/100)^3\) = ?

ganeshie8 (ganeshie8):

yw ! :)